MODS1 Commissioning Run 4: UTC 2011 January 12

Observer(s): R. Pogge (OSU), M. Pedani (LBTO)
Telescope Operator: S. Allanson (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Little (LBTO)
Instrument Support: T. Shih (LBTO)

Summary

Last night one of the wheel bearings on one of the Bogies that the building uses to turn (Bogie B) died with an awful sound. The enclosure has been frozen in place since last night. Inspection by Bob Meeks and company so far has found two damaged outer bearings, the one that died on Bogie B and another on Bogie A. Inspection of all of the bogie wheels is not yet complete.

As such there is no observing tonight. Too bad, it is cirrusy but might have been useful.

Using the night for software updates, maintenance, and calibration in zenith-pointing mode.

Weather was bands of cirrus until about 0930 when the sky become overcast with heavy high clouds (no stars visible except occasional thinning). Temperatures ranged from 0 to -1°C, Humidity was between 20 and 40%. Winds were 3-8m/sec out of the W. When the wind was below 10m/s we used the vent doors to try to better equilibrate the outside air and mirror temperatures. The mirrors are still quite far from good thermal equilibrium.

Preparations

Instrument woken up w/o problems. Total time from reboot to ready-to-observe (i.e., test images acquired, IMCS lock, archive and TCS links verified) was under 15 minutes.

Test bias images look normal.

Details

All times are UTC unless otherwise noted.

-- RichardPogge - 12 Jan 2011
Topic revision: r4 - 12 Jan 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback